Collins Group
The Collins Group, a leading producer of custom automobile accessories, has hired you to estimate the firm's weighted average cost of capital. The balance sheet and some other information are provided below.

Assets
Current assets

$  38,000,000

Net plant, property, and equipment

  101,000,000

Total assets

$139,000,000

Liabilities and Equity
Accounts payable

$  10,000,000

Accruals

      9,000,000

Current liabilities

$  19,000,000

Long-term debt (40,000 bonds, $1,000 par value)

    40,000,000

Total liabilities

$  59,000,000

Common stock (10,000,000 shares)

30,000,000

Retained earnings

    50,000,000

Total shareholders' equity

    80,000,000

Total liabilities and shareholders' equity

$139,000,000

The stock is currently selling for $15.25 per share, and its noncallable $1,000 par value, 20-year, 7.25% bonds with semiannual payments are selling for $875.00. The beta is 1.25, the yield on a 6-month Treasury bill is 3.50%, and the yield on a 20-year Treasury bond is 5.50%. The required return on the stock market is 11.50%, but the market has had an average annual return of 14.50% during the past 5 years. The firm's tax rate is 25%.

1. Refer to the data for the Collins Group. Which of the following is the best estimate for the weight of debt for use in calculating the firm's WACC?

a.

22.69%

b.

21.61%

c.

18.67%

d.

19.60%

e.

20.58%

2. Refer to the data for the Collins Group. Based on the CAPM, what is the firm's cost of common stock?

a.

13.00%

b.

12.35%

c.

11.73%

d.

13.65%

e.

11.15%
Answer #1

Solution:

Question 1 .

We need to find the weight of debt, we have to find the market value of debt and market value of equity.

Market value of Debt = Current price of bond * Number of bonds = 875*40,000 = 35,000,000

Market value of equity = Current share price * Number of share = 15.25 *10,000,000 = 152,500,000

Weight of debt = Market value of debt / ( Market value of debt + Market value of Equity ) = 35,000,000 /(35,000,000 + 152,500,000) = 18.67%

Option C is correct

Question 2.

Risk free rate = 5.5% ( Since bond has maturity of 20 years hence we have chosen this time horizon for risk free rate)

Using CAPM,

Cost of common stock = Risk free rate + Beta * (Market return - Risk free rate ) = 5.50% + 1.25 * (11.50% - 5.50%) = 5.5% + 7.5% = 13%

Correct option is A ) 13%